05-30-2009
May 30, 2009
ISS On-Orbit Status 05/30/09

All ISS systems continue to function nominally, except those noted previously or below. Saturday – first full day for six-member Expedition 20, with the crew of CDR Gennady Padalka (Russia), FE-1 Michael Barratt (USA), FE-2 Koichi Wakata (Japan), FE-3 Roman Romanenko (Russia), FE-4 Robert Thirsk (Canada), FE-5 Frank DeWinne (Belgium).

The FE-2 continued his third session of sleep logging for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. [To monitor the crewmember’s sleep/wake patterns and light exposure, Koichi wears a special Actiwatch device which measures the light levels encountered by them as well as his patterns of sleep and activity throughout the Expedition and uses the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

Padalka & Romanenko performed the periodic maintenance of the active Russian BMP Harmful Impurities Removal System by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The process will be terminated tonight at ~3:30pm EDT before crew sleep, followed tomorrow by Bed #2 regeneration. (Last time done: 5/10-5/11). [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle is normally done every 20 days.]

The full crew performed the regular weekly task of thorough station cleaning, including COL (Columbus Orbital Laboratory) and Kibo. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, the CDR & FE-3 conducted regular maintenance inspection & cleaning on fan screens in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9), plus dust filter replacement in the FGB.

Romanenko also performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, while Padalka did the periodic maintenance of the Russian POTOK-150MK (150 micron) air purification subsystems in the SM & FGB by cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting the POTOK in automatic mode. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Continuing preparations in the DC1 (Docking Compartment) for the Orlan suited dry-run on 6/3 (Wednesday) and the EVA-22 spacewalk on 6/5 (Friday), CDR Padalka & FE-1 Barratt checked out the Orlan “Korona” communication links, which include VHF/voice and biomedical electrode belt and telemetry hookups via the BSS interface units (to be switched after egress to the Tranzit-B wireless in-suit radio telemetry system). [The tests of the VHF/voice & biomedical electrode belt and telemetry hookups for vital signs and equipment monitoring involved a ground specialist tagup over RGS (Russian Groundsite) via VHF at 3:55-4:15am EDT.]

Afterwards, as part of the regular physical fitness check prior to an Orlan spacewalk, Gennady & Mike undertook the Russian MO-5 MedOps protocol of cardiovascular assessment during graded physical load on the VELO cycle ergometer, assisting each other in turn as CMO (Crew Medical Officer). [The assessment uses the Gamma-1 ECG equipment with biomed harness, skin electrodes and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer's instrumentation panels. Measurements were telemetered down via VHF to RGS (Russian Groundsite) during two comm windows (3:55am & 6:45am). For the graded-load exercise, the subject works the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each. Data output involves a kinetocardiogram, rheoplethysmogram, rheoencephalogram and a temporal pulsogram.]

Floating in the Habitation Module of the newly arrived Soyuz TMA-15/19S spacecraft, Roman installed the LKT local temperature sensor commutator (TA251M1B) of the BITS2-12 onboard telemetry system and its ROM/read-only memory unit (PZU TA765B), both kept in storage from an earlier Soyuz.

Padalka performed the frequent status check on the Russian BIO-5 Rasteniya-2 ("Plants-2") experiment, verifying proper operation of the BU Control Unit and MIS-LADA Module fans (testing their air flow by hand) plus today, as a discretionary task list item, monitoring seedling growth, humidity measurements, moistening of the substrate if necessary and topping off the water tank if ~20-25% of the total amount (4 liters) remains. [Rasteniya-2 researches growth and development of plants under spaceflight conditions in the LADA-15 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP).]

The full crew completed the formal 2-hr Safety Briefing, reviewing in detail procedures and escape routes in case of an emergency. [The Safety Handover included safety-related items such as emergency actions (egress path to the two Soyuz vehicles, procedures, annunciations, isolation equipment, safety equipment, depressurization and leak repair, fire, toxic release), visiting vehicles docking/undocking, evacuation vehicles, crew life support system status, computers, communications, medical equipment, IVA hazards (e.g., sharp edges, protrusions, touch temperatures), stowage, and current hardware status.]

At ~9:00am EDT, the crew conducted the regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

After Wakata again set up video coverage of the TVIS (Treadmill with Vibration Isolation & Stabilization) worksite, the CDR continued troubleshooting the critical exercise device, first deactivating the three remaining TVIS stabilizers, then conducting another exercise run on the treadmill, using an alternate method (see Note below). Koichi later downlinked the video & audio for ground inspection.

Roman & Frank each had 15 min set aside for familiarizing themselves with the physical exercise procedures onboard and particularly the current usage of the TVIS.

The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS (CDR), ARED advanced resistive exercise device (FE-1, FE2) and VELO cycle ergometer with bungee cord load trainer (CDR, FE-1, FE-2). [On the CEVIS, the actual loads remain slightly lower than the commanded loads, but this was expected. A manual correction of the pertinent calibration coefficient via the control panel touch screen will be done at a later time when the new value has been determined.]

Later, Wakata transferred the exercise data file to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

The three newcomers had another PMC (Private Medical Conference) via S- & Ku-band audio/video, their second, Frank DeWinne at ~7:40am, Roman Romanenko at ~2:00pm, Bob Thirsk at ~2:50pm EDT.

A new addition on Mike Barratt’s voluntary “job jar” task list was some more troubleshooting of the AgCam (Agricultural Camera) system, after his successful fix of AgCam laptop issues on 5/23 during his VolSci (Voluntary Science) activity.[When attempting to take the first image earlier this week, another anomaly, unrelated to the first one, became apparent, involving use of cables and components. For this weekend, Mike’s voluntary task (changing to a backup cable and data card plus re-seating several other cables) was to enable AgCam to complete science prior to window shutter closing on 6/9 (will remain closed until 2J/A departure).]

On Gennady’s “time permitting” discretionary task list, another session of the GFI-8 "Uragan" (hurricane) earth-imaging program called for NIKON D2X digital camera photography (with 800mm telelens) of high priority targets such as fires on the Volga-Akhtyubinsk alluvial plain and in the Volga river delta, floods in the Volga-Akhtyubinsk alluvial plain and Volga river delta, Krasnaya Polyana and its mountain ski resort, the Laganaki plateau, etc.

TVIS Update: The TVIS treadmill continues to be evaluated and is not cleared yet for normal usage. When Koichi Wakata yesterday conducted the TVIS exercise checkout session with stabilizers in the passive mode, they unexpectedly became active at a gyroscope speed of about 2000 rpm, causing again the anomalous overdriving oscillations. When returning to ~1800 rpm and below, the stabilizers switched back to passive mode. Ground analysis continues, and until this phenomenon is explained and the problem is resolved, an uplinked alternate method for running on the TVIS with stabilizers disabled was used by Gennady today.

FPMU SDTO Activities: Yesterday, ground controllers began an extended SDTO (Station Demonstration Test Objective) with the FMPU (Floating Potential Measurement Unit) to evaluate ISS vehicle charging with this system. [The final test associated with this SDTO, it will provide a complete picture of solar array-driven plasma charging of ISS for the entire complement of ISS SAs (solar arrays) and also address the effectiveness of ISS plasma shock hazard mitigation techniques. SAs are being positioned in order to generate sufficient charges of high potential; the arrays are vertical (90 deg) and BGAs (Beta Gimbal Assemblies) are positioned in forward (“ram“) or rearward (“wake”)-facing directions. Both Russian & European Langmuir probes have been activated in support of the test. The JEM PROX system is also activated during this SDTO to evaluate the performance of the RGPS (Relative Global Positioning System) and verification of the multi-path environment, both of which are critical to HTV (H-II Transfer Vehicle) success. The SDTO will continue until 6/5 (EVA-22). Due to high momentum at the beginning of the FPMU testing, with SARJs (Solar Alpha Rotary Joints) in directed position of 90 deg, desaturation firings by Russian thrusters have been enabled (last time done: 2004) to desaturate the U.S. CMGs (Control Moment Gyroscopes) as required.]

Onboard Water Balance: With the crew doubled in size, work is underway to re-assess the onboard water situation and establish a new balance for the USOS (US Segment) with its WRS (Water Recovery System), which now has more water going in and more coming out. [To provide empirical data for re-determining the balance, Mike, Koichi, Frank & Bob were asked to maximize use of the water from the WRS for now.]

Weekly Science Update (Expedition Nineteen -- Week 8)

3-D SPACE: Ongoing.

AgCam (Agricultural Camera): “Mike, the AgCam team was very happy with your successful laptop exchange during your Voluntary Science activity on Saturday, and were all geared up to get good imagery this week. However on Wednesday, 5/27, when attempting to take their first image, another anomaly, unrelated to the first one, became apparent. The AgCam team believes a short troubleshooting crew activity to change a cable would do the trick, and a task list activity to do that is under development.”

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): BCAT-4 is complete. BCAT-5: June launch on 2J/A.

BIOLAB (ESA): No report.

Biological Rhythms (JAXA): Ongoing. Two runs are for Barratt, two are planned for De Winne.

BISE (CSA, Bodies in the Space Environment): CSA/Canada staff and the PI team look forward to see Bob Thirske and Frank DeWinne perform the upcoming BISE sessions.

BISPHOSPHONATES: “Koichi, thanks for completing your Alendronate ingestion activity. Your next pill ingestion is scheduled for Monday, 6/1.”

CARD (Long Term Microgravity: A Model for Investigating Mechanisms of Heart Disease, ESA): On 5/27, the data from the CARD blood pressure measurements recorded during the CARD Day#1 protocol on 5/20, were downlinked as part of other ground commanded EPM (European Physiology Module) activities. Data is now being analyzed by CARD science team.

CARDIOCOG-2: Complete.

CBEF (JAXA Cell Biology Experiment Facility): Troubleshooting continues. “The CBEF dehumidification was completed. Humidity reached to certain level to be able to start data acquisition of the CO2 sensor which has sensitive for humidity. Fan Rotations have also been finished by Koichi. Thank you for providing more effort than the scheduled task. The results show that the fan of the temp control unit in the micro gravity compartment is completely failed. The ground engineering team is now investigating how to repair this function. The first repair task will be prepared and need to be performed before running the Space Seed experiment, currently scheduled during the 17A stage.”

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Mike, thanks for completing your second CCISS session. Thanks for skipping your lunch and repeating the Baro and 24-HR Heart Rate Study. You are the first subject to complete the FD60 session. We are investigating the Holter Monitor 2 anomaly.”

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): During last week’s six test point attempts (three methanol, three heptane), two methanol burns were successful but only one of the heptane burn attempts was partially successful. The first methane droplet drifted into the igniter, which immediately stopped the burn. The first two heptane attempts did not burn while the third heptane droplet hung on the dispenser needle and did not have a true microgravity characteristic during its burn. The heptane test point will be re-run at TBD date. The team is still assessing future operations with the given stowage configuration around the CIR rack.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Completed.

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): “Thank you Mike for having successfully performed the exchange of the ESEM Board. During this activity, the PCDF instrument was powered down for 2hrs in total, which is short enough from a science point of view. In order to minimize the temperature fluctuations for the PCDF reactors during the planned power down period, Columbus Moderate Temperature Loop (MTL) set-point was increased from +17degC to +19degC. While the ESEM showed good H&S data from the re-activation of EDR, some anomalies have been encountered during PCDF restart (see dedicated PCDF paragraph). The EDR status is fully nominal since 5/27.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

ENose (Electronic Nose): Operating.

EPM (European Physiology Module): Underway.

EPO (Educational Payload Operations): Reserve.

EPO J-Astro Report (JAXA): “Koichi’s mission is coming close to the end with two reports remaining. Thanks!”

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): Complete.

EPO Moon Score (JAXA): Planned.

EPO Try Zero-G (JAXA): “The last session will be scheduled as Voluntary Science on the June 6,7 weekend. The former two sessions were released on the JAXA website. It is a very good performance. Thank you for your continuous effort.”

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): “Additional photo was very interesting. Thank you. Please keep photo in proprietary folder when you download them.”

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): EuTEF platform and its instruments have been running nominally during this reporting period.-- DEBIE-2: New script was started on 5/27. The power cycle work-around has been performed on 5/26 and 5/27.-- EXPOSE: On-going science acquisition. On 5/13, the valves of the Experimental Trays have been closed for the NH3 venting event. It is planned to keep those closed for about 2 weeks to avoid any contamination to the samples;-- FIPEX: The last EOP script has been nominally performed until 5/25. New script was started on 05/27 (GMT147); -- MEDET: the instrument has been manually restarted on 5/19 and since then is acquiring science;-- PLEGPAY: On 5/21, PLEGPAY Langmuir Probe measurements have been resumed (after ~8.5mths of stand-down due to safety concerns). Preliminary analysis of the measurements shows that the Langmuir Probe is in good shape. Science measurements have been resumed between 5/26 and 5/28. PLEGPAY instrument will be de-activated one day prior to 19S docking.

FACET (JAXA): “75% of experiments have finished. Data show that Microgravity condition in JEM is really good for FACET experiment. We planned two more weeks to run the experiment and all will finish just before the 2J/A arrival if everything goes well.”

FSL (Fluid Science Laboratory): No report.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): Complete.

HQPC (JAXA): To be launched by 34P.

ICE CRYSTAL (JAXA): Complete.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

Integrated Immune: Complete.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

MAXI (JAXA): Ongoing.

Micro-G Clay (JAXA EPO): Complete.

MMA (Microgravity Measurement Apparatus): First attempt using ELT (Experiment Laptop Terminal) with MMA for micro-G data acquisition was successfully finished after a laptop exchange from MLT (MMA Laptop Terminal) to ELT. We successfully acquired background microgravity data in the JEM using the modified configuration. The MLT hard drive was removed by Koichi (thank you for quick task) and expected to be retrieved by 2J/A for micro-G data salvage on the ground. New hard drive installation will be scheduled after 2J/A which will carry a new drive unit.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: Ongoing.

PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): Over the last weekend, one additional science run has been performed with the EP3 reactor from 5/22 to 5/24. On 5/25, science acquisition started with the last of the 4 PCDF reactors, i.e. EP2. This last reactor is dedicated to study the phenomena of protein depletion zone apparition during crystal growth phase. Initial measurements had to be interrupted in order to perform the EDR ESEM board exchange on 5/27. Both PCDF Electronics Unit (EU) and PCDF Process Unit (PU) were powered off for a total duration of 2hrs during this EDR ESEM board exchange. After this crew activity, EDR got re-activated, followed by PCDF-EU and PCDF-PU power on sequence. Very rapidly an anomaly occurred during the PCDF-PU boot-up, as it started to rapidly cool down in an unexpected way (science impact TBC as soon as new images are available on ground). Ground team immediately performed a power cycle of the PCDF-PU, and normal temperature control of the PCDF-PU was then recovered, but leading to different operative mode of the PU (i.e. Stand-Alone Mode), which did not allow to resume the science acquisition. PCDF was kept in this configuration overnight, with a higher temperature set-point of the Columbus Moderate Temperature Loop (MTL) at +19degC (instead of usual +17degC). On 5/28, ESA engineering experts gave their recommendation to proceed with a full and clean power cycling of both PCDF-EU and PCDF-PU. After ground commanding, PCDF-PU was put back in a nominal configuration and the science acquisition could be restarted. The EP4 reactor still contains some nice crystals to be returned by 2J/A. Next weeks will be devoted to grow additional crystals in EP2 reactor.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

POLCA/GRAVIGEN (ESA): Complete.

RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Koichi, you are currently completing your second week of sleep logging this increment. Mike, thanks for completing your second week of sleep logging. The data were sent to the PI for analysis.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): The current Sun visibility window has started on 5/22, one day earlier than predicted. SOLAR platform is in Sun Tracking Mode since then. The present Sun observation window is predicted to end on 6/5. – SOLSPEC: on-going science acquisition, with regular calibration cycles with operational lamps and using a work-around calibration procedure for the UV spectrum; -- SOLACES: on-going science acquisition.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): Planned.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): Through 5/26, the ground has received a total of 9,487 of ISS CEO imagery for review and cataloguing. [“Normal review of your imagery has resumed with a sizeable backlog. No imagery with current dates, since we resumed daily target lists this week, have been received.We are pleased to confirm your acquisition of views of the following targets: Key Largo, Florida; Madeira Island, Portugal; Ascension Island, Atlantic Ocean; and Galapagos Islands, Ecuador. Last weekend, your striking image of a large circular thinning of the ice in southern Lake Baikal, Siberia was published on NASA/GSFC’s Earth Observatory website. Astronauts have photographed similar features in the lake just two other times since 1985. Their origin remains undetermined. This weekend another of your images, an interesting view of remote, solar evaporations salt ponds in the Atacama Desert of northern Chile will also be posted. The green and yellow colors of the ponds of a large nitrate mining operation stand in stark contrast to the brown alluvial fans of this hyperarid region. These are both examples of alert documentary photography of rarely seen features on our Earth. Kudos for your efforts!”]

CEO (Crew Earth Observation) photo targets uplinked for today were Irkutsk, Russia (Irkutsk is one of the largest cities in Siberia. The city is located to the north of the southwestern end of Lake Baikal. Looking to the left of track for the city – from the Lake, following the Irkutsk River northwest; when the channel narrows significantly, there is Irkutsk. Overlapping frames of the metropolitan area were requested), Paris, France (weather was predicted to be clear over Paris. Overlapping, nadir-viewing frames as ISS approached, passed over and departed the metropolitan area were requested), Bonn, Germany (clear weather conditions are also predicted for Bonn. This city, the former capital of West Germany, is located along the Rhine River is the western part of Germany. The larger city of Köln (Cologne) is located downstream to the northwest. Overlapping frames, taken slightly to the left of track, provide a rural-urban-rural transect across the city), and Merida, Mexico (Yucatan Peninsula) (some patchy clouds may have been present over Merida. This nadir-viewing overpass provided an opportunity to collect a rural-urban-rural transect across the city area).

CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this noon, 1:05pm EDT [= epoch])
Mean altitude -- 349.6 km
Apogee height – 356.1 km
Perigee height -- 343.2 km
Period -- 91.53 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0009585
Solar Beta Angle -- -5.9 deg (magnitude peaking)
Orbits per 24-hr. day -- 15.73
Mean altitude loss in the last 24 hours -- 128 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 60320

Significant Events Ahead (all dates Eastern Time, some changes possible!):
06/03/09 -- Orlan Suited Dry-Run (training)
06/05/09 -- Russian EVA-22
06/10/09 -- Russian EVA-23
06/13/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD (7:12am)
06/29/09 -- STS-127/Endeavour/2J/A landing (12:18am EDT, KSC)
07/17/09 – Progress M-02M/33P undock & deorbit
07/20/09 -- Soyuz TMA-14/18S relocation (from SM aft to DC1)
07/24/09 -- Progress 34P launch
07/26/09 -- Progress 34P docking (SM aft)
08/06/09 -- STS-128/Discovery/17A – MPLM (P), LMC
09/01/09 -- H-IIB (JAXA HTV-1) launch – tentative
09/07/09 -- H-IIB (JAXA HTV-1) berth
09/30/09 -- Soyuz TMA-16/20S launch
10/02/09 -- Soyuz TMA-16/20S docking (SM aft, until MRM-2 w/new port)
10/08/09 -- H-IIB (JAXA HTV-1) unberth
10/11/09 – Soyuz TMA-14/18S undock
10/15/09 -- Progress 35P launch
11/10/09 -- 5R/MRM-2 (Russian Mini Research Module 2) on Proton -- tentative
11/12/09 -- STS-129/Atlantis/ULF3 - ELC1, ELC2
12/07/09 -- Soyuz TMA-17/21S launch
12/26/09 -- Progress 36P launch
02/03/10 -- Progress 37P launch
02/XX/10 -- STS-130/Endeavour/20A – Node-3 + Cupola -- tentative
02/11/10 -- STS-131/Atlantis/19A – MPLM(P), LMC -- tentative
03/05/10 -- Progress 38P launch
04/02/10 -- Soyuz TMA-18/22S launch
04/08/10 -- STS-132/Discovery/ULF4 – ICC-VLD, MRM-1 -- tentative
04/30/10 -- Progress 39P launch
05/30/10 -- Soyuz TMA-19/23S launch
06/30/10 -- Progress 40P launch
07/29/10 -- STS-133/Endeavour/ULF5 – ELC3, ELC4 -- tentative
07/30/10 -- Progress 41P launch
09/30/10 -- Soyuz TMA-20/24S launch
10/30/10 -- Progress 42P launch
11/??/10 -- ATV2 – Ariane 5 (ESA)
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton